what is enthalphy change when 952 mg of hydrogen gas reacts completely with carbon monoxide


1
Expert's answer
2021-04-16T04:59:36-0400

As a result of the interaction of carbon monoxide with hydrogen methane gas and water are formed. Reaction equation is: CO + 3H2 = CH4 + H2O. The reaction should be carried out with heating in the temperature range 150-200 °C and with the participation of a catalyst - nickel.

Then, ∆H reaction = ∆H products - ∆H reactants

∆H reaction = - 74.848 - 285.84 + 110.523 = -250.165 kJ/mol

n(H2) = 0.952/2 = 0.476 mole

Enthalpy change = 0.476×(-250.165) / 3 = -39.69 kJ

This is one of the options for solving this problem.

In the case of the reaction, it will be higher than 230 - 400 °C, instead of nickel in the catalyst there will be copper (II) oxide or chromium (III) oxide, and also play under pressure, then instead of methane, you can get methyl alcohol (methanol). Reaction equation is:

CO + 2H2 = CH3OH

Then, ∆H reaction = - 238.64 + 110.523 = -128.117 kJ/mol

Enthalpy change = 0.476×(-128.117) / 2 = -30.69 kJ